Please Help!
Consider the following circuit 10.0 Ω R2 5.00 Ω 2.00 w R1-6.00 Ω If the current flowing through the 2.002 resistor is 2.00 A, and the current flowing through the 5.00 Ω resistor is 0.44 A, determine the following values: The current flowing through R1, R2, the equivalent resistance, and &
Answer #1

Here, if the current in 5 Ohm is 0.44 A

for the current in 10 Ohm

10 * I1 = 5 * 0.44

I1 = 0.22 A

current in R2 = 0.22 + 0.44 = 0.66 A

Now, current in R1 = 2 - 0.66 = 1.34 A

NOw, for as the potential across the upper and lower branch is same

let the upper branch have resistance R3

R3 * 0.66 = 1.34 * 6

R3 = 12.2 Ohm

for the net resistance

equivalent resistance = 12.2 * 6/(12.2 + 6) + 2 = 6.02 Ohm

the equivalent resistance is 6.02 Ohm

for the emf

E = 2 * 6.02 = 12.04 V

the emf E is 12.04 V

current in R1 is 1.34 A

current in R2 is 0.66 A